# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

### v3

@@ Abstract
We compare two time series analysis methods, the Discrete Fourier Transform (DFT) and our Discrete Chi-square Method (DCM). DCM is designed for detecting signal(-s) superimposed on an unknown trend. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational statistical well-posed solution for this problem. The backbone of DCM is the Gauß-Markov theorem that the least squares fit is the best unbiased estimator for linear regression models. DCM can not fail because this simple time series analysis method computes a massive number of linear least squares fits. Hence, the data spacing, even or uneven, is irrelevant. We use the Fisher-test to identify the best DCM model from all alternative tested DCM models. This best model must also pass our Predictivity-test. Our analyses of simulated complex data sets expose the weaknesses of DFT and the efficiency of DCM. The DFT and DCM are frequency-domain parametric methods. There are many other similar parametric methods. Just like DFT, those other methods suffer from their own particular application limitations. The list of those limitations is long. DCM suffers from none of those limitations. The performance of DCM depends only on the data set size and accuracy. DCM is an ideal forecasting method because the data set time span $(\Delta T)$ is irrelevant. It does not matter how long $(\Delta T)$ and/or complex the data set is because DCM will inevitably detect the signal(-s) and the trend when the data set size $(n)$ and/or accuracy $(\sigma)$ become adequate.
@@
Computational statistics, Ill-posed problems, Non-linear time series analysis models, Discrete Chi-square Method, Discrete Fourier Transform
@@ 1 Introduction
[24] formulated the least squares method. [13] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the least-squares method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [14]. The extended Gauß-Markov theorem states that the least squares method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [34]. The Gauß-Markov theorem is the backbone of our DCM. This time series analysis method performs a massive number of linear least squares fits to find the best model for the data.
@@
There are numerous techniques for solving the ill-posed problem of detecting one signal superimposed on a trend, especially for evenly spaced data [7, 28, 2]. Kay and Marple [22] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data. They discussed how the time span of data $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another. The spectral estimating techniques fail if the data are non-stationary. Therefore, trends changing the sample mean and/or variance must be removed before applying spectral estimation [26]. Ghaderpour et al. [15] presented different techniques for reducing spectral leakage. They also discussed techniques for detecting many signals in unevenly spaced data, if these signals are pure sines.
However, no published solution seems to exist for the more complex data sets, where the signal periods are longer than $\Delta T$ and/or these signals are not pure sines. We will show that DCM can find a well-posed computational statistical solution even for such complex data sets. This method can detect an unknown number of periodic signals superimposed on an unknown aperiodic trend. The noise level can be unknown because the performance of DCM depends only on the sample size $(n)$ and data accuracy $(\sigma)$. The time span $(\Delta T)$ is irrelevant. The data spacing, even or uneven, is also irrelevant.
In this study, we apply one published DFT version [19] to analyse artificial simulated unevenly spaced data having equal weights. The performance of this DFT version is compared to the performance of our DCM.
We present the DCM formulation in Section 2.1. Data sets are simulated using the DCM model (Sect. 2.2). The four DFT application limitations are discussed in Section 2.3 (Equations 27-30). DCM and DFT time series analysis of the simulated data sets shows that DCM does not suffer from the DFT application limitations (Sects. 3.1-3.7). We then present the identification of the best DCM model for the data (Sect. 3.8), the DCM significance estimates (Sect. 3.9) and the DCM predictions (Sect. 3.10). Finally, we discuss our results (Sect. 4).
@@ 2 Methods
The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n). The time span of data, which is also called the data window, is $\Delta T=t_{n}-t_{1}$. The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$. The mean and the standard deviation of all $y_{i}$ values are denoted with $m$ and $s$.
@@ 2.1 Discrete Chi-square method (DCM)
[21] formulated DCM. The model of this method is
@@
Due to this group, all free parameters are not eliminated from all partial derivatives $\partial g/\partial\beta_{i}$. This makes the $g(t)$ model non-linear. If the $\bm{\beta}_{I}$ frequencies have constant values, the multipliers $2\pi jf_{i}$ in Equation 2.1 become constants. In this case, the model becomes linear because the partial derivatives $\partial g/\partial\beta_{i}$ no longer contain any free parameters. The solution for the second group of remaining free parameters
@@
Let us assume that we search for periods between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. The non-linear $g(t)$ model becomes linear, if the tested $\bm{\beta}_{I}$ frequencies are fixed to any constant values. The sum $h(t)$ of signals $h_{1}(t,f_{1})$, $h_{2}(t,f_{2})$ … and $h_{K_{1},f_{K_{1}}}(t)$ does not depend on the order in which these signals are added. For example, the two signal $K_{1}=2$ model symmetry is $h(t)=h_{1}(t,f_{1})+h_{2}(t,f_{2})=h_{1}(t,f_{2})+h_{2}(t,f_{1})$. Both $(f_{1},f_{2})$ and $(f_{2},f_{1})$ combinations give the same value for the $g(t)$ model. Since this symmetry applies to any $K_{1}$ number of signals, we compute the linear $g(t)$ models only for all tested frequency combinations
@@
For every tested $\bm{\beta}_{I}=[f_{1},f_{2},...,f_{K_{1}}]$ frequency combination, the DCM test statistic
$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{R/n}$ (15)
$z$ $=$ $z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{\chi^{2}/n}$ (16)
is computed from a linear model least squares fit. Equation 15 or Equation 16 is applied for unknown or known $\sigma_{i}$ errors, respectively. The value of $z$ is unique.
In the preliminary long search, we test an evenly spaced grid of $n_{\mathrm{L}}$ frequencies between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. This search gives the best frequency candidates $f_{\mathrm{1,mid}},...f_{\mathrm{K_{1},mid}}$.
@@
where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively. The even or uneven data spacing is irrelevant because the linear least squares fit result for every tested frequency combination does not depend on this spacing.
@@
In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 19) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$. These one-dimensional periodogram slices (Equation 20) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 15 and 16). This kind of visualisation is important in time series analysis [29].
@@
  • $P_{i}=1/f_{i}=$ Period
  • $A_{i}=$ Peak to peak amplitude
  • $t_{\mathrm{i,min,1}}=$ Deeper primary minimum
  • $t_{\mathrm{i,min,2}}=$ Secondary minimum (if present)
  • $t_{\mathrm{i,max,1}}=$ Higher primary maximum
  • $t_{\mathrm{i,max,2}}=$ Secondary maximum (if present)
@@
would be practically impossible because the number of DCM model free parameters is large, especially if the data contains many signals. We use the computational statistical bootstrap technique [10] to solve these analytical problems. This approach gives the DCM model parameter error estimates, as well as the $\sigma_{g(t)}$ model error estimates (Equation 22). In our bootstrap, the tested frequencies are the same as in the short search. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 12). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create numerous artificial bootstrap data sets
$\bm{y}^{*}=\bm{g}+\bm{\epsilon}^{*}.$ (23)
Each $\bm{y}^{*}$ sample gives one estimate for every DCM model parameter value and model $g(t)$ value. For each particular model parameter and each particular $g(t)$, the bootstrap error estimate is the standard deviation of all estimates obtained from all $\bm{y}^{*}$ samples.
@@
Intersecting frequencies occur when the signal frequencies in the data are very close to each other. We give the following example of how this instability can arise in the two signal model. If the frequency $f_{1}$ approaches the frequency $f_{2}$, both $h_{1}(t)$ and $h_{2}(t)$ signals become essentially one and the same signal. The least squares fit fails because it makes no sense to model the same signal twice.
Dispersing amplitudes instability can occur, if the two signal frequencies are too close to each other. The least squares fit finds a model, where two high amplitude signals nearly cancel out each other. The low amplitude signal, the sum of these two high amplitude signals, fits to the data.
@@
2.2 Simulated DCM model data samples
We use seven different $g(t)$ models to simulate 21 artificial complex data sets (Sections 3.1-3.7). The time interval of all simulated data sets is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [22] is always

### v4

# Discrete Chi-Square Method can model and forecast complex time series

@@ Abstract
We show how intensive, large and accurate time series can allow us to see through time. Many phenomena have aperiodic and periodic components. An ideal time series analysis method would detect such trend and signal(-s) combinations. The widely-used Discrete Fourier Transform (DFT) and other frequency-domain parametric time series analysis methods have many application limitations constraining the trend and signal(-s) detection. We show that none of those limitations constrains our Discrete Chi-square Method (DCM) which can detect signal(-s) superimposed on an unknown trend. Our simulated time series analyses ascertain the revolutionary Window Dimension Effect (WD-effect): “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.” The simulations also expose the DFT’s weaknesses and the DCM’s efficiency. The DCM’s backbone is the Gauß-Markov theorem that the Least Squares (LS) is the best unbiased estimator for linear regression models. DCM can not fail because this simple method is based on the computation of a massive number of linear model LS fits. The Fisher-test gives the signal significance estimates and identifies the best DCM model from all alternative tested DCM models. The analytical solution for the non-linear DCM model is an ill-posed problem. We present a computational well-posed solution. The DCM can forecast complex time series. The best DCM model must be correct if it passes our Forecast-test. Our DCM is ideal for forecasting because its WD-effect spearhead is robust against short sample windows and complex time series.
@@
Time series analysis, Non-linear and non-stationary processes, Ill-posed problems, Computational statistics, Forecasting, Discrete Chi-square Method, Discrete Fourier Transform
@@ 1 Introduction
Cheng et al. [7] stated, “Forecasting the evolution of complex systems is noted as one of the 10 grand challenges of modern science.” We will show that the DCM can detect the signal(-s) and the trend in complex time series, and can also forecast the time series evolution. Our DCM is based on the LS method, which was originally formulated by [28]. [14] connected this method to the principles of probability and the normal distribution. A few years later, he showed that the LS method gives the best unbiased estimates for the free parameters of linear models, if the zero mean data errors are equal, normally distributed and uncorrelated [15]. The extended Gauß-Markov theorem states that the LS method gives the best estimates for the free parameters of linear models, even if the data errors do not pass all the above-mentioned criteria [43]. The Gauß-Markov theorem is the backbone of our DCM. This time series analysis method performs a massive number of LS fits to find the best model for the data.
@@
Our DCM is a frequency-domain parametric time series analysis method. Here, we discuss shortly such methods in the order of increasing model complexity. Numerous methods can be applied if the time series is stationary (constant mean and variance). The Fast Fourier Transform for evenly spaced data [8] and the DFT for unevenly spaced data [30, 36] are the most widely-used methods. The FFT and DFT assume that the correct model for the time series is a pure sine. The spectral estimating techniques fail if the data are non-stationary. Therefore, trends changing the sample mean and/or variance must be removed before applying spectral estimation [33]. There are numerous methods for detecting one signal superimposed on a trend, especially for evenly spaced data [6, 37, 2]. Kay and Marple [25] compared many spectral estimating techniques that can detect more than one signal from evenly spaced data. They discussed how the sample window $(\Delta T)$ limits the frequency resolution for detecting many signals and causes leakage that shifts power from one spectral peak to another. Exactly correct frequency values may not be detected because the leakage can shift the periodogram peaks. Ghaderpour et al. [17] presented different techniques for reducing spectral leakage. They also discussed techniques for detecting many signals in unevenly spaced data, if these signals are pure sines. There are more complex modelling problems. For example, the signals are not necessarily pure sines or the sample window $\Delta T$ is shorter than the signal period(-s).
Reliable forecasting is the ultimate test for any time series analysis method [7]. Forecasting can work for linear and stationary processes, but it is challenging for non-linear or non-stationary ones. Other forecasting challenges are overfitting and forecast error estimation [32, 27, 16, 34].
Our DCM is a frequency-domain parametric time series analysis method. Such methods have their own particular limitations that constrain their applications. Here is our list of those application limitations (AL).
  1. 1.
    Data errors (level of noise) are unknown.
  2. 2.
    Data error information is not utilised.
  3. 3.
    Data must be evenly spaced.
  4. 4.
    Model parameter errors are unknown.
  5. 5.
    Model and forecast errors are unknown.
  6. 6.
    Sample window is shorter than signal period(s).
  7. 7.
    Presence and shape of trend are unknown.
  8. 8.
    Sample window causes leakage.
  9. 9.
    Leakage weakens frequency resolution.
  10. 10.
    Signal shapes are not pure sines.
  11. 11.
    Number of signals is unknown.
  12. 12.
    Correct model alternative is unknown.
  13. 13.
    Signal significances are unknown.
  14. 14.
    Model solution is ill-posed.
  15. 15.
    Complex non-linear model forecasts fail.
This study proceeds through following stages. We present the DCM formulation (Section 2.1). The WD-effect is discussed (Section 2.2). Complex time series are simulated using the DCM model (Sect. 2.3). The DFT and its ALs are presented (Section 2.4). We compare our DCM to the renowned DFT. The time series analysis of simulated data sets shows that the ALs of DFT do not constrain DCM (Sections 3.1-3.8). Then, we present the identification of the best DCM model for the data (Section 3.9), the DCM significance estimates (Section 3.10), the well-posed computational DCM model solution (Section 3.11) and the DCM forecasts (Section 3.12). Finally, we discuss our results (Section 4).
@@ 2 Methods
The observations are $y_{i}=y(t_{i})\pm\sigma(t_{i})$, where $t_{i}$ are the observing times and $\sigma_{i}$ are the errors (i=1,2,,n)(i=1,2,...,n). The sample window is $\Delta T=t_{n}-t_{1}$. The mid point is $t_{\mathrm{mid}}=t_{1}+\Delta T/2$. The mean and the standard deviation of all $y_{i}$ values are denoted with $m$ and $s$.
@@ 2.1 Discrete Chi-square method (DCM)
[23] introduced the DCM which is a frequency-domain parametric time series analysis method. The primary objective of this current paper is to show that the DCM outperforms all other similar methods.
The DCM model is
@@
Due to this group, all free parameters are not eliminated from all partial derivatives $\partial g/\partial\beta_{i}$. This makes the $g(t)$ model non-linear. If the $\bm{\beta}_{I}$ frequencies have constant values, the multipliers $2\pi jf_{i}$ in Equation 2.1 become constants. In this case, the model becomes linear because the partial derivatives $\partial g/\partial\beta_{i}$ no longer contain any free parameters. The LS solution for the second group of remaining free parameters
@@
Let us assume that we search for periods between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. The non-linear $g(t)$ model becomes linear, if the tested $\bm{\beta}_{I}$ frequencies are fixed to any constant values. The sum $h(t)$ of signals $h_{1}(t,f_{1})$, $h_{2}(t,f_{2})$ … and $h_{K_{1},f_{K_{1}}}(t)$ does not depend on the order in which these signals are added. For example, the two signal $K_{1}=2$ model symmetry is
$h(t)$ $=$ $h_{1}(t,f_{1})+h_{2}(t,f_{2})$
$=$ $h_{1}(t,f_{2})+h_{2}(t,f_{1})$
Both $(f_{1},f_{2})$ and $(f_{2},f_{1})$ combinations give the same value for the $g(t)$ model. Since this symmetry applies to any $K_{1}$ number of signals, we compute the linear $g(t)$ models only for all tested frequency combinations
@@
This frequency space symmetry idea reduces CPU consumption dramatically. For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space. This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube). The search for the best frequency combination from $4!=24$ different non-linear iterations would be a pointless exercise (see Equation 22).
@@
For every tested $\bm{\beta}_{I}=[f_{1},f_{2},...,f_{K_{1}}]$ frequency combination, the LS fit gives the DCM test statistic
$z$ $=z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{R/n},$ unknown $\sigma_{i}$ (16)
$z$ $=z(f_{1},f_{2},...,f_{K_{1}})=\sqrt{\chi^{2}/n},$ known $\sigma_{i}$ (17)
computed for a linear model. The value of $z$ is unique. The errors can be unknown in Equation 16 (AL1). For known errors, DCM uses this information in Equation 17 (AL2). In the preliminary long search, we test an evenly spaced grid of $n_{\mathrm{L}}$ frequencies between $f_{\mathrm{min}}$ and $f_{\mathrm{max}}$. This search gives the best frequency candidates $f_{\mathrm{1,mid}},...f_{\mathrm{K_{1},mid}}$.
@@
where $n_{f}=n_{L}$ and $n_{f}=n_{S}$, respectively. The even or uneven data spacing is irrelevant because the LS fit result for every tested frequency combination does not depend on this spacing (AL3).
@@
In the above-mentioned $K_{1}=2$ map, the slice $z_{1}(f_{1})$ would represent the height $z$ at the location $(f_{1},f_{\mathrm{2,best}})$ when moving along the straight constant line $f_{2}=f_{\mathrm{2,best}}$ that crosses the global minimum $z_{\mathrm{min}}$ (Equation 20) at the coordinate point $(f_{\mathrm{1,best}},f_{\mathrm{2,best}})$. These one-dimensional periodogram slices (Equation 21) allow us "to see inside” the multi-dimensional structure of the DCM test statistic $z$ (Equations 16 and 17). This kind of visualisation is important in time series analysis [38].
@@
$P_{i}$ $=$ $1/f_{i}=\text{Period}$ (23)
$A_{i}$ $=$ Peak to peak amplitude (24)
$t_{\mathrm{i,min,1}}$ $=$ Deeper primary minimum (25)
$t_{\mathrm{i,min,2}}$ $=$ Secondary minimum (if present) (26)
$t_{\mathrm{i,max,1}}$ $=$ Higher primary maximum (27)
$t_{\mathrm{i,max,2}}$ $=$ Secondary maximum (if present) (28)
@@
would be practically impossible because the number of DCM model free parameters is large, especially if the data contains many signals.
We use the computational statistical bootstrap technique [11] to solve these analytical problems. In our bootstrap, the tested frequencies are the same as in the short search. We select a random sample $\bm{\epsilon}^{*}$ from the residuals $\bm{\epsilon}$ of the DCM model (Equation 13). Any $\epsilon_{i}$ value can be chosen to the $\bm{\epsilon}^{*}$ sample as many times as the random selection happens to favour it. We create $J=1,2,...,n_{\mathrm{B}}$ residual random samples $\bm{\epsilon}^{*}_{J}$. Every $\bm{\epsilon}^{*}_{J}$ sample gives one artificial bootstrap data set
$\bm{y}^{*}_{J}=\bm{g}+\bm{\epsilon}^{*}_{J}.$ (31)
Each $\bm{y}^{*}_{J}$ sample gives one free parameter estimate $\bm{\beta}_{J}$. This bootstrap procedure gives $n_{\mathrm{B}}$ free parameter estimates $\bm{\beta}_{J}$. The standard deviations for all $n_{\mathrm{B}}$ estimates for $P_{i}=1/f_{i}$ and all $M_{k}$ give the errors of these model parameters. We use each $\bm{\beta}_{J}$ to compute the model for a dense grid of time points. This gives us $n_{\mathrm{B}}$ numerical $A_{i}$ $t_{\mathrm{i,min,1}}$, $t_{\mathrm{i,min,2}}$, $t_{\mathrm{i,max,1}}$ and $t_{\mathrm{i,max,2}}$ estimates. The standard deviation of those $n_{\mathrm{B}}$ estimates is the error of these model parameters. The $\bm{\beta}_{J}$ values give $n_{\mathrm{B}}$ estimates for $g(t)$ at any time $t$. The standard deviation of these $g(t)$ estimates gives the error $\sigma_{g(t)}$ of Equation 30. Note that these $\bm{\beta}_{J}$ values can also be used to compute the errors for $h(t),h_{i}(t),p(t)$ and $p_{k}(t)$. Our computational statistical bootstrap approach gives not only the errors estimates for the DCM model parameters (AL4), but also the model error inside $\Delta T$ and the forecast error outside $\Delta T$ (AL5: Equation 30).
@@
Intersecting frequencies occur when the signal frequencies in the data are very close to each other. We give the following example of how this instability can arise in the two signal model. If the frequency $f_{1}$ approaches the frequency $f_{2}$, both $h_{1}(t)$ and $h_{2}(t)$ signals become essentially one and the same signal. The LS fit fails because it makes no sense to model the same signal twice.
Dispersing amplitudes instability can occur, if the two signal frequencies are too close to each other. The LS fit finds a model, where two high amplitude signals nearly cancel out each other. The low amplitude signal, the sum of these two high amplitude signals, fits to the data.
@@
The DCM model (Equation 1) is more sophisticated than the models of our former time series analysis methods, the Three Stage Period Search [24, TSPA, Equation 1] and the Continuous Period Search [29, CPS, Equation 3]. TSPA and CPS can detect only one signal $(K_{1}=1)$ from stationary time series $(K_{3}=0)$. TSPA can detect pure sine and double wave signals ($K_{2}=1$ or 2). The extension of TSPA, the CPS method, tests three alternatives: $K_{2}=1$ or 2, or no signal at all. The DCM model is more sophisticated because it can have any arbitrary $K_{1}$, $K_{2}$ and $K_{3}$ combination. The DCM sum $g(t)$ =h(t)+p(t)=h(t)+p(t) of arbitrary periodic and aperiodic functions represents a universal model because the innumerable $K_{1}$, $K_{2}$ and $K_{3}$ combinations allow unlimited complexity. Even more complex signal shape $(K_{2})$ combinations could be used. For example, these shapes could be $K_{2,i=1}=3$ (3rd harmonic), $K_{2,i=2}=1$ (pure sine) and $K_{2,i=3}=2$ (double wave), where if $K_{2,i}$ denotes the shape of $i$:th signal.
2.2 WD-effect: Spearhead of DCM
In this section, we discuss what causes the WD-effect defined in our abstract:
  • “For any sample window $\Delta T$, DCM inevitably detects the correct $p(t)$ trend and $h(t)$ signal(-s) when the sample size $n$ and/or data accuracy $\sigma$ increase.”
The consequences of this effect are also discussed.
The $t_{i}$, $y_{i}$ and $\sigma_{i}$ data are inside the rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i}+\sigma_{i})-{\mathrm{min}}(y_{i}-\sigma_{i})]$. The LS fit results do not depend on $\Delta T$. If the measuring time intervals $\delta t_{i}$ for each observation $y_{i}$ fulfil $\delta t_{i}\ll\Delta T$, the $R$ and $\chi^{2}$ values obtained from all LS (Equation 19) depend only on $y_{i}$ changes in $\Delta Y$ direction, but not on the $t_{i}$ changes in $\Delta T$ direction. For fixed $y_{i}$ and $\sigma_{i}$, the residuals $\epsilon_{i}$ determine the $R$ and $\chi^{2}$ values. These residuals measure only the $\Delta y$ direction. Hence, the periodogram $z$ values (Equation 16 or 17) do not depend on $\Delta T$. If $\Delta T$ decreases, the $R$ and $\chi^{2}$ estimates can measure the $g_{i}$ model details inside the $\Delta T\times\Delta y$ rectangle only if $n$ increases and/or $\sigma$ decreases. Better data reveal these model details.
Plenty of concrete examples will confirm that the WD-effect is real (Sections 3.1-3.7). Our DCM can detect signals(-s) when the sample window $\Delta T$ is shorter than the period $P$ value(-s). The DCM surpasses AL6. The data window $\Delta T$ length is irrelevant. This means that the DCM can model an infinitesimal time series, as well as forecast its future and past, if the sample size $(n)$ and/or the data accuracy $(\sigma)$ are sufficient. That is a revolutionary achievement.
2.3 Simulated DCM model time series samples
We use seven different $g(t)$ models to simulate 21 artificial complex time series (Sections 3.1-3.7). The sample window of all simulated time series is $\Delta T=1.$ The $n^{\star}$ simulated time points $t_{i}^{\star}$ are drawn from a random uniform distribution $U(0,\Delta T,n^{\star})$. The first and last time point values are then modified to $t_{1}^{\star}=0$ and $t_{n}^{\star}=\Delta T$. Hence, the distance between independent frequencies [25] is always